OpenStudy (anonymous):

what is the solution to 8 - 2x ( greater then or equal to ) -4 ? My question is do you add 4 to both sides ?

OpenStudy (amistre64):

you can .... it wont hurt in the long run

OpenStudy (amistre64):

i would add 4 and add 2x to each side

OpenStudy (anonymous):

okay.thanks.

OpenStudy (anonymous):

could you show me how you'd do that ?

OpenStudy (anonymous):

\[\LARGE 8-2x \geq -4\] \[\LARGE -2x \geq -4-8\] \[\LARGE -2x \geq -12\quad \quad \quad |\cdot (-1)\] \[\LARGE 2x \leq 12\quad \quad \quad | :(2)\] \[\LARGE x \leq 6\]